Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
rolland - 9:37 Monday 17 November 2025 (68172) Print this report
PCal switched off (following storm and timing issue)

I received DMS alerts about the Pcal power this morning at 8h16 and again 8h26 UTC. The Pcal looked off from the VIM plots; there were errors messages in the Acl processes about missing/delayed samples. And Alain confirmed that there has been a strom and that the timing has probably been cut.

At 8h30 UTC, I have stopped the laser Pcal (open loop and disable the power from Acl PCal_fast, and switched off the pump diodes of the laser from PCal_WE,NE_Ampli).  We will switch them on when the timing will be back.

Comments to this report:
rolland - 8:47 Tuesday 18 November 2025 (68179) Print this report

The PCal were switched on yesterday around 13h30 (NE) and 14h30 (WE) UTC.  (entry from yesterday but not posted).

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×